ashley is 3/4 as old as her brother. if the sum of their ages is 28 years, how old is each?
Answer by Timnewman(323)

Let his brother's age be x
Ashley's age=3x/4
If the sum of their ages is 28,
Then,
x+(3x/4)=28
multiply out by 4
4x+3x=112
7x=112
x=16yrs
Ashley;s brother;s age is 16yrs
Ashley's age=(3*16)/4
=48/4
=12yrs
HOPE THIS HELPS?
GOOD LUCK!